OpenStudy (anonymous):

lim t>7 F(x) (1/t)−(1/7)/(t-7)

OpenStudy (anonymous):

\[\lim_{t \rightarrow 7} \frac{\frac{1}{t} - \frac{1}{7}}{t-7} \]?

OpenStudy (anonymous):

\(\lim_{t \rightarrow 7} \frac{\frac{1}{t} - \frac{1}{7}}{t-7} \) \(\lim_{t \rightarrow 7} \frac{\frac{7-t}{7t}}{t-7} \) \(\lim_{t \rightarrow 7} \frac{\frac{t-7}{-7t}}{t-7} \) \(\lim_{t \rightarrow 7} \frac{1}{-7t} \) do you need further help?

OpenStudy (anonymous):

Thanks

OpenStudy (anonymous):

you're welcome :)